An express and local train leave Gray's Lake at 3PM and head for Chicago 50 miles away. The express travels twice as fast as the local and arrive 1 hour ahead of it. Find the speed of each train.
Answer by mangopeeler07(462)

2x=express speed
x=local speed
50=miles

50/2x=(50/x)-1

Common denominator=2x
50/2x=100/2x-2x/2x

Eliminate the denominator
50=100-2x

Subtract 100 from both sides
-50=-2x

Divide by -2
25=x

2x=express speed
x=local speed

Local train: 25mph
Express train: 50mph